OpenStudy (anonymous):

Solve for x: 1/5 x + 2 = 8. 50 30 40 10

thomaster (thomaster):

\(\Large\sf \dfrac{1}{5x+2}=8\) ?

OpenStudy (anonymous):



thomaster (thomaster):

Oke \(\Large\sf \dfrac{1}{5}x+2=8\) First get the 2 to the right

OpenStudy (anonymous):

First you want to get rid of the fraction. You do that by times each side by 5

OpenStudy (anonymous):